OpenStudy (mama2dylan):

Write an exponential function. A population of 120,000 grows 1.2% per year for 15 years.

jimthompson5910 (jim_thompson5910):

The general form is \[\Large F = P(1+r)^{t}\] F = final population (after t years) P = initial population r = growth rate (if negative, then you have exponential decay) t = time in years

jimthompson5910 (jim_thompson5910):

also, keep in mind that 1.2% = 0.012

OpenStudy (mama2dylan):

Thank you. Can you write the function out using the figures I provided @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You will plug in these values P = 120000 r = 0.012 t = 15 and when you use a calculator, you can compute the value of F (the final population after t = 15 years)

jimthompson5910 (jim_thompson5910):

so this basically \[\Large F = P(1+r)^{t}\] \[\Large F = 120000(1+0.012)^{15}\] \[\Large F =\ ???\]

jimthompson5910 (jim_thompson5910):

it seems odd how they want a function, but they say t = 15. So maybe they want you to not replace t with 15 and just leave it as \[\Large F(t) = P(1+r)^{t}\] \[\Large F(t) = 120000(1+0.012)^{t}\] \[\Large F(t) = 120000(1.012)^{t}\] but I'm not sure

OpenStudy (perl):

i think its saying that the exponential function's domain is from t =0 to t=15

OpenStudy (perl):

since it is growing 1.2% each year

OpenStudy (perl):

over an interval of 15 years